# Value-Free Policy Optimization via Reward Partitioning

### v2

@@ Abstract
Single-trajectory reinforcement learning (RL) methods aim to optimize policies from datasets consisting of (prompt, response, reward) triplets, where scalar rewards are directly available. This supervision format is highly practical, as it mirrors real-world human feedback, such as thumbs-up/down signals, and avoids the need for structured preference annotations. In contrast, pairwise preference-based methods like Direct Preference Optimization (DPO) rely on datasets with both preferred and dispreferred responses, which are harder to construct and less natural to collect. Among single-trajectory approaches, Direct Reward Optimization (DRO) has shown strong empirical performance due to its simplicity and stability. However, DRO requires approximating a value function, which introduces several limitations: high off-policy variance, coupling between policy and value learning, and a lack of absolute supervision on the policy itself. We introduce Reward Partitioning Optimization (RPO), a new method that resolves these limitations by removing the need to model the value function. Instead, RPO normalizes observed rewards using a partitioning approach estimated directly from data. This leads to a straightforward supervised learning objective on the policy, with no auxiliary models and no joint optimization. RPO provides direct and stable supervision on the policy, making it robust and easy to implement in practice.
We validate RPO on scalar-feedback language modeling tasks using Flan-T5 encoder-decoder models. Our results demonstrate that RPO outperforms existing single-trajectory baselines such as DRO and Kahneman-Tversky Optimization (KTO). These findings confirm that RPO is a simple, effective, and theoretically grounded method for single-trajectory policy optimization. Our code is publicly available at https://github.com/b-faye/rpo.
@@ I Introduction
The dominant approach for aligning large language models (LLMs), whether through reinforcement learning from human feedback (RLHF) [1, 2, 3] or direct preference optimization, is to learn from preference or reward data [4, 5, 6, 7]. In RLHF, models are first trained via supervised fine-tuning (SFT) [8], followed by learning a reward model (RM) from human-labeled preference pairs (e.g., two responses to the same prompt ranked by quality), and finally aligned with reinforcement learning (e.g., PPO [9], TRPO [10]) to maximize the RM’s score. This pipeline is central to the training of models such as InstructGPT [11], GPT-4 [12], and Claude 3 [13].
While effective, RLHF presents key limitations: collecting preference data and training reward models is costly, and the RL step is sample-inefficient and brittle. To reduce reliance on human annotation, recent work has proposed Reinforcement Learning from AI Feedback (RLAIF), where feedback is generated by models rather than humans. RLAIF includes strategies such as distilling AI feedback into a reward model [14], using an LLM directly as a reward function [15, 16], or leveraging self-evaluation [17, 18]. These techniques offer scalability but still depend on large models as evaluators, reintroducing cost and inference complexity.
A parallel line of work avoids reward models and RL altogether, by directly optimizing policies from preference data via supervised learning. These methods—such as DPO [4], GPO [5], and SLiC [19]—contrast response pairs to a prompt and define a preference-based loss. This strategy enables simple and stable training, and has been adopted in models like LLaMA 3 [20], Qwen 2 [21], and Nemotron-4 [22].
However, collecting high-quality pairwise preference data is increasingly difficult. As LLMs improve, distinguishing between good responses becomes subtler, demanding more qualified annotators and driving up labeling costs. This makes scaling preference-based methods impractical at the frontier.
To address this, recent work proposes single-trajectory learning, where data consists of a prompt, a response, and a scalar reward (e.g., thumbs-up/down). This format is simpler and more natural for annotation. Methods such as KTO [7] and DRO [6] leverage this setup to train policies directly from reward-labeled completions.
Among them, DRO has shown strong empirical results by regressing the optimality condition from KL-regularized RL. However, DRO suffers from several limitations: it requires learning a value function from off-policy data, the joint optimization over policy and value is unstable, and the policy is only supervised relatively to a reference model, not directly by reward.
@@
We introduce Reward Partitioning Optimization (RPO), a new method for single-trajectory policy optimization. RPO eliminates the need to estimate a value function by computing an empirical normalization over observed rewards. This enables a direct supervised loss on the policy, using only prompt-response-reward triplets and a reference model.
RPO is simple to implement, avoids instability from value approximation, and provides absolute reward supervision. It is fully offline-compatible and scalable to large datasets. Our contributions are:
  • We analyze theoretical and practical limitations of DRO for single-trajectory learning.
@@
  • We propose RPO, a new method that replaces value estimation with empirical reward normalization.
  • We provide a formal derivation showing RPO approximates the optimal policy structure.
@@
  • We validate RPO on instruction-tuning tasks using T5 encoder-decoder models [23], demonstrating strong performance against DRO and KTO baselines.
@@ II Background
In this section, we review existing paradigms for aligning large language models (LLMs) with human feedback, focusing on three main categories: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style). We examine their assumptions, limitations, and motivation for a more general framework that can directly leverage scalar reward data.
@@ II-A Reinforcement Learning from Human Feedback (RLHF)
RLHF has become the dominant paradigm for aligning large language models (LLMs) with human preferences. It proceeds in three main stages:
@@
  • Reward Model (RM) Training: Given a dataset $\mathcal{D}$ of pairwise preferences $(x,y^{\text{win}},y^{\text{lose}})$—where $x$ is a prompt, and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred response—the goal is to learn a scalar reward function $r_{\phi}(x,y)$ that reflects human preferences. This is typically modeled using a Bradley-Terry [24] formulation:
@@
    $\footnotesize\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))].\$ (2)
@@
    where $\tau>0$ is a regularization coefficient that balances reward maximization and policy deviation. Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [9] are used to perform the optimization.
Despite its success, RLHF is resource-intensive and challenging to implement. Training the reward model demands significant human supervision and is susceptible to out-of-distribution generalization errors that can mislead the policy. Furthermore, reinforcement learning algorithms add substantial computational overhead and are sensitive to hyperparameter tuning. These drawbacks have spurred interest in alternative alignment strategies that bypass explicit reinforcement learning.
@@ II-B Direct Preference Optimization (DPO-style)
To address the high cost and complexity of full reinforcement learning, a family of recent approaches—such as DPO [4], IPO [25], and SLiC [19]—propose optimizing policies directly from preference data without learning a separate reward model or relying on reinforcement learning algorithms like PPO.
Given a dataset of pairwise preferences $\mathcal{D}=\{(x,y^{\text{win}},y^{\text{lose}})\}_{i=1}^{N}$, where $x$ is a prompt and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred completion, these methods define an objective that compares the log-likelihoods of the winning and losing responses under the learned policy $\pi_{\theta}$, relative to a fixed reference policy $\pi_{\text{ref}}$. The general form of the loss is:
$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\left[f\left(\beta\left(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\right)\right)\right],$ (4)
where $\beta>0$ is a temperature parameter and $f$ is a shaping function that varies by method:
  • DPO uses the logistic loss: $f(z)=\log(1+e^{-z})$;
  • SLiC uses a hinge loss: $f(z)=\max(0,1-z)$;
  • IPO uses a squared loss centered at $z=1$: $f(z)=(z-1)^{2}$.
This loss encourages the policy $\pi_{\theta}$ to assign higher probability to preferred responses than to dispreferred ones, while penalizing large deviations from the reference policy $\pi_{\text{ref}}$. In essence, it minimizes a divergence between the preference-consistent ranking and the policy-induced ranking, thereby aligning the model’s behavior with human preferences without requiring a learned reward function or reinforcement learning procedure.
These methods can be interpreted as KL-regularized preference optimization, where the model is trained to favor preferred completions relative to the reference while remaining close to $\pi_{\text{ref}}$. Crucially, they bypass the need for an explicit reward model and reinforcement learning, leading to simpler and more stable training.
However, their reliance on manually curated preference datasets poses a practical limitation. In real-world deployment, feedback often comes in the form of scalar or binary reward signals, i.e., as $(x,y,r)$ tuples rather than preferences—motivating alternative approaches capable of leveraging such logged data directly.
@@ II-C Direct Reward Optimization (DRO-style)
DRO-style refers to a class of methods that directly optimize a policy from scalar rewards rather than pairwise preferences. Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$, where $x$ is a prompt, $y$ a model response, and $r\in\{-1,+1\}$ a scalar reward indicating human feedback, DRO-style methods aim to maximize the utility of desirable responses while minimizing that of undesirable ones—without relying on an explicit reward model or reinforcement learning algorithm.
A representative example of this approach is the KTO [7] method, which draws inspiration from prospect theory [26] to model human preferences with loss aversion and risk sensitivity.
Let $\pi_{\theta}$ denote the current policy and $\pi_{\mathrm{ref}}$ a fixed reference policy. Define the log-likelihood ratio between the two as:
$\footnotesize r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}.$ (5)
Next, define a reference point $z_{0}$ as the expected log-likelihood ratio over sampled responses:
$\footnotesize z_{0}=\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot\mid x)}\left[\log\frac{\pi_{\theta}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]=\mathrm{KL}(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)).$ (6)
To model asymmetric treatment of gains and losses, KTO introduces the following components:
  • A temperature parameter $\beta>0$ to control the sharpness of value sensitivity;
  • A logistic function $\sigma(z)=\frac{1}{1+e^{-z}}$ for smooth transition between reward regions;
  • Distinct loss aversion coefficients $\lambda_{D}$ (for desirable responses) and $\lambda_{U}$ (for undesirable ones).
The value function $V(x,y)$ is then defined as:
$\footnotesize V(x,y)=\begin{cases}\lambda_{D}\cdot\sigma\left(\beta\left[r_{\theta}(x,y)-z_{0}\right]\right),&\text{if }r=+1,\\ \lambda_{U}\cdot\sigma\left(\beta\left[z_{0}-r_{\theta}(x,y)\right]\right),&\text{if }r=-1.\end{cases}$ (7)
Finally, the KTO objective minimizes the following expected loss over the data distribution:
$\footnotesize\mathcal{L}_{\mathrm{KTO}}(\theta)=\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\lambda_{r}-v(x,y)\right],$ (8)
where $\lambda_{r}=\lambda_{D}$ if $r=+1$ and $\lambda_{U}$ if $r=-1$.
This formulation allows KTO to smoothly reward desirable generations and penalize undesirable ones, while incorporating human-like asymmetries in preference treatment. Unlike binary cross-entropy or advantage-weighted likelihood objectives, KTO provides a principled, prospect-theoretic loss tailored to scalar reward data.
KTO models human risk aversion but relies on strong assumptions and optimizes only the policy $\pi_{\theta}$, ignoring a separate value function $V(x)$. In contrast, DRO [6] jointly optimizes both $\pi$ and $V$, offering a more principled and general framework.
Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$ where $r$ is a scalar reward, DRO formulates policy learning as a KL-regularized reward maximization problem:
$\footnotesize\pi^{*}=\arg\max_{\pi}\;\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}\left[r(x,y)-\tau\cdot\mathrm{KL}\left(\pi(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x)\right)\right],$ (9)
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\tau>0$ is a temperature hyperparameter controlling the strength of regularization.
The optimal policy $\pi^{*}$ can be expressed in closed form as:
$\footnotesize\pi^{*}(y|x)=\pi_{\mathrm{ref}}(y|x)\cdot\exp\left(\frac{1}{\tau}r(x,y)-\frac{1}{\tau}V^{*}(x)\right),$ (10)
where $V^{*}(x)$ is the soft value function, defined by
$\footnotesize V^{*}(x)=\tau\log\mathbb{E}_{y\sim\pi_{\mathrm{ref}}(\cdot|x)}\left[\exp\left(\frac{1}{\tau}r(x,y)\right)\right].$ (11)
Rearranging, this identity links reward, value, and policy:
$\footnotesize r(x,y)=V^{*}(x)+\tau\cdot\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (12)
To approximate the optimal $(\pi^{*},V^{*})$, DRO minimizes the squared Bellman residual:
$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2}\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\left(r-V_{\phi}(x)-\tau\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right)^{2}\right],$ (13)
where $\pi_{\theta}$ and $V_{\phi}$ denote parameterized policy and value functions respectively.
In practice, this loss is optimized jointly over parameters $\theta$ and $\phi$ using the observed data samples $\{(x_{i},y_{i},r_{i})\}_{i=1}^{n}$ as:
^$\footnotesize\hat{\mathcal{L}}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2N}\sum_{i=1}^{N}\left(r_{i}-V_{\phi}(x_{i})-\tau\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}\right)^{2}.$ (14)
This principled framework optimizes policies directly from scalar rewards by jointly learning a value function and applying KL regularization for stability. However, this joint optimization often suffers from instability, especially when the data distribution differs from the policy, as value estimation errors can bias training. To address these issues, we propose RPO, which retains DRO’s strengths but removes the need to estimate a value function, simplifying training and enhancing robustness.

### v3

@@ Abstract
Single-trajectory reinforcement learning (RL) trains policies from datasets of (prompt, response, reward) triplets, where scalar rewards provide direct supervision. This format aligns with real-world human feedback and avoids the cost of pairwise preference annotations required by methods like Direct Preference Optimization (DPO). Among single-trajectory methods, Direct Reward Optimization (DRO) is effective but depends on value-function approximation, leading to high variance, policy–value coupling, and weak direct supervision.
We propose Reward Partitioning Optimization (RPO), which eliminates value modeling by normalizing rewards via data-driven partitioning. This yields a purely supervised policy objective, free of auxiliary models and joint optimization. RPO thus offers stable, direct supervision and practical simplicity.
Experiments on scalar-feedback language modeling with Flan-T5 show that RPO consistently outperforms single-trajectory baselines including DRO and Kahneman-Tversky Optimization (KTO), establishing RPO as a simple, robust, and theoretically grounded approach to single-trajectory policy optimization.
@@ I Introduction
Aligning large language models (LLMs) typically relies on preference or reward data, via reinforcement learning from human feedback (RLHF) [NEURIPS2023_23e6f78b, dong2024rlhf, ahmadian] or direct preference optimization [rafailov2023direct, tang2024, richemond2024offline, ethayarajh2024]. In RLHF, models are first fine-tuned via supervised learning (SFT) [zhang2023instruction], then a reward model (RM) is trained from human-labeled preference pairs, and finally aligned with reinforcement learning (e.g., PPO [schulman2017proximal], TRPO [schulman2015trust]) to maximize RM scores. This pipeline underlies InstructGPT [ouyang2022training], GPT-4 [achiam2023gpt], and Claude 3 [anthropic2024claude].
RLHF is effective but costly: collecting preference data, training reward models, and RL steps are sample-inefficient and brittle. Reinforcement Learning from AI Feedback (RLAIF) [xu2025, du2023guiding, yuan2023rrhf, wang2025self, yuan2024self] replaces humans with AI evaluators, improving scalability but still depending on large models and introducing inference overhead.
An alternative avoids reward models and RL entirely, directly optimizing policies via supervised learning on preference pairs [rafailov2023direct, tang2024, zhao2023slic]. While simple and stable, high-quality pairwise data is increasingly hard to obtain as LLMs improve, making scaling impractical [grattafiori2024llama, yang2025qwen3, adler2024nemotron].
Single-trajectory learning simplifies supervision: datasets contain (prompt, response, reward) triplets (e.g., thumbs-up/down). Methods like KTO [ethayarajh2024] and DRO [richemond2024offline] train policies directly from these scalar-labeled completions. DRO performs well but requires learning a value function from off-policy data, suffers unstable joint optimization, and supervises the policy only relative to a reference model.
@@
We propose Reward Partitioning Optimization (RPO), which eliminates value estimation by normalizing observed rewards empirically. This yields a direct supervised loss on the policy using only prompt-response-reward triplets and a reference model. RPO is simple, stable, offline-compatible, and scales to large datasets. Our contributions are:
  • Analysis of DRO’s theoretical and practical limitations for single-trajectory learning.
@@
  • Introduction of RPO, replacing value estimation with empirical reward normalization.
@@
  • Validation on instruction-tuning tasks with T5 encoder-decoder models [raffel2020exploring], demonstrating strong performance against DRO and KTO baselines.
@@ II Background
In this section, we review major approaches for aligning large language models (LLMs) with human feedback: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style). We highlight their assumptions and limitations, motivating a broader framework that directly leverages scalar reward data.
@@ II-A Reinforcement Learning from Human Feedback (RLHF)
RLHF has become the dominant paradigm for aligning LLMs with human preferences. It proceeds in three main stages:
@@
Reward Model (RM) Training: Given a dataset $\mathcal{D}$ of pairwise preferences $(x,y^{\text{win}},y^{\text{lose}})$—where $x$ is a prompt, and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred response—the goal is to learn a scalar reward function $r_{\phi}(x,y)$ that reflects human preferences. This is typically modeled using a Bradley-Terry [bradley1952rank] formulation:
@@
$\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))]$
@@
where $\tau>0$ is a regularization coefficient that balances reward maximization and policy deviation. Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal] are used to perform the optimization.
Despite its success, RLHF is costly and difficult to implement: reward models require heavy human supervision and risk out-of-distribution errors, while reinforcement learning adds computational overhead and sensitivity to tuning. These challenges motivate alignment methods that avoid explicit RL.
@@ II-B Direct Preference Optimization (DPO-style)
To reduce the cost of full RL, recent methods such as DPO [rafailov2023direct], IPO [azar2024general], and SLiC [zhao2023slic] optimize policies directly from preference pairs without training a reward model or running PPO.
Starting from the KL-regularized objective in Equation 3 and explicitly enforcing the normalization constraint $\sum_{y}\pi_{\theta}(y|x)=1$, the Lagrangian yields a closed-form expression for the optimal policy:
$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y)\right)}{\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y^{\prime})\right)}.$ (4)
Given a dataset of preference pairs
$\mathcal{D}=\{(x_{i},y_{i}^{\text{win}},y_{i}^{\text{lose}})\}_{i=1}^{N},\quad y_{i}^{\text{win}}\succ y_{i}^{\text{lose}},$
we can express the reward difference directly from Equation 4 as
$\footnotesize\frac{r_{\phi}(x_{i},y_{i}^{\text{win}})-r_{\phi}(x_{i},y_{i}^{\text{lose}})}{\tau}=\log\frac{\pi_{\theta}^{\ast}(y_{i}^{\text{win}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{win}}|x_{i})}{\pi_{\theta}^{\ast}(y_{i}^{\text{lose}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{lose}}|x_{i})}.$ (5)
Minimizing a surrogate loss based on this advantage leads to the general preference-based objective:
$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\Bigg[f\Bigg(\beta\Big(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\Big)\Bigg)\Bigg],$
where $\beta>0$ is a temperature, and $f$ depends on the method:
  • DPO: $f(z)=-\log\sigma(z)$
  • SLiC: $f(z)=\max(0,1-z)$ (hinge loss)
  • IPO: $f(z)=(z-1)^{2}$ (squared loss)
This KL-regularized loss pushes $\pi_{\theta}$ to rank preferred completions above dispreferred ones while staying close to $\pi_{\text{ref}}$, enabling stable training without explicit reward models or RL. However, reliance on curated preference data limits scalability, since real-world feedback often comes as scalar rewards $(x,y,r)$-motivating methods that can exploit such signals directly.
@@ II-C Direct Reward Optimization (DRO-style)
DRO-style methods optimize policies directly from scalar rewards rather than preference pairs. Given a dataset $\mathcal{D}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$, where $r_{i}$ denotes scalar feedback, these methods maximize the utility of desirable responses without explicit reinforcement learning or reward model training.
A notable example is KTO [ethayarajh2024], which draws inspiration from prospect theory [tversky1992advances] to model loss aversion and risk sensitivity. In contrast, DRO [richemond2024offline] starts from the KL-regularized optimal policy in Equation 4:
$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\tfrac{1}{\tau}r(x,y)\right)}{\exp\left(\tfrac{1}{\tau}V^{\ast}(x)\right)},$ (7)
where
$V^{\ast}(x)=\tau\log\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\!\left(\tfrac{1}{\tau}r(x,y^{\prime})\right)$
is the optimal soft value function.
Rearranging Equation 7 yields the consistency condition:
$\footnotesize r(x,y)=V^{\ast}(x)+\tau\log\frac{\pi_{\theta}^{\ast}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (8)
DRO approximates $(\pi^{\ast},V^{\ast})$ by jointly learning a policy $\pi_{\theta}$ and a value function $V_{\phi}$, parameterized by a neural network with parameters $\phi$, through minimization of the Bellman residual:
$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\tfrac{1}{2}\,\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\Big[\big(r-V_{\phi}(x)-\tau\log\tfrac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\big)^{2}\Big].$ (9)
This objective requires joint optimization of the policy and value network, which can introduce instability due to off-policy value estimation errors. Our proposed RPO eliminates the need for value function learning, simplifying optimization while retaining the advantages of direct reward-based policy learning.